Given the following​ after-tax cash flow on a new toy for​ Tyler's Toys, find the​ project's payback​ period, NPV, and IRR. The appropriate discount rate for the project is 9​%. If the cutoff period is 6 years for major​ projects, determine whether management will accept or reject the project under the three different decision models.

Initial cash​ outflow: ​$12,800,000
Years one through four cash​ inflow: ​$3,200,000 each year
Year five cash​ outflow: ​$1,280,000
Years six through eight cash​ inflow: ​$548,667 each year

What is the payback period for the new toy at​ Tyler's Toys?
nothing years  ​(Round to two decimal​ places.)
Under the payback​ period, this project would be
accepted
rejected

What is the NPV for the new toy at​ Tyler's Toys?
Under the NPV​ rule, this project would be
accepted
rejected

What is the IRR for the new toy at​ Tyler's Toys?
Under the IRR​ rule, this project would be
rejected
accepted
Answer #1

1.
Payback=12800000/3200000
=4.00

2.
Under payback, accept

3.
NPV=-12800000+3200000/1.09+3200000/1.09^2+3200000/1.09^3+3200000/1.09^4-1280000/1.09^5+548667/1.09^6+548667/1.09^7+548667/1.09^8=-2362159.258

4.
Under NPV rule reject

5.
IRR=IRR({-12800000;3200000;3200000;3200000;3200000;-1280000;548667;548667;548667})=1.011%

6.
Under IRR rule reject